# Linear constellations in primes with arithmetic restrictions

### v2

@@ 9.4. W-trick
The following lemma is a simple reduction completely analogous to the deduction of [13, Theorem 5.1] from [13, Theorem 5.2].
@@
We need the following estimate in our application of the $W$-trick in Lemma 9.9. Its proof is, essentially, contained in the proof of [13, Lemma 1.3].
@@
so that $\psi_{i}(Wn+c)=W\psi_{i,c}(n)+b_{i}(c)$. Note that if $\gcd(W,\psi_{i}(c))>1$, then $\gcd(\psi_{i}(Wn+c),W)>1$ for any $Wn+c\in X$. As $W\ll_{\mathbf{a}}(\log N)^{2/3}$ and $\psi_{i}\geq N^{8/10}$ on $X$, this shows that $\psi_{i}(Wn+c)$ can not be prime if $N$ is large enough. Hence, identifying $[W]^{s}$ with $(\mathbb{Z}/W\mathbb{Z})^{s}$ in the obvious way, we see that
@@
In [13, §4], the main theorem of [13] is first reduced to [13, Theorem 4.1], removing the archimedean factor, and then futher to [13, Theorem 4.5], extending $\Psi$ to a system in normal form.
@@
As we have a pseudorandom majorant already, the deduction of Theorem 1.4 from Proposition 2.1 follows the same steps as in §§9.39.5 with the obvious modifications. For example, we rely on the results concerning $\tau_{\mathbf{K},\mathbf{C},\Psi}(q)$ established in §3.2 instead of the analogous properties of $\sigma_{\mathbf{a},\Psi}(q)$ established in §3.3. Again, the situation is somewhat simpler, as the function $w(\cdot)$ is fixed once and for all and does not depend on any of the other parameters. The argument is essentially the same as in [13, §§4-7].
@@ 11.1. One linear equation in primes with prescribed Artin symbols
Here we show that the main term Kane’s result [23, Theorem 2] for $t\geq 3$ matches up with a special case or our Theorem 1.2. We may formulate Kane’s asymptotic as follows: for $t\geq 3$ and $\mathbf{K},\mathbf{C}$ as in Theorem 1.4, $\mathbf{c}=(c_{1},\ldots,c_{t})\in\mathbb{Z}^{t}$ with $\gcd(c_{1},\ldots,c_{t})=1$ and $N,M\in\mathbb{N}$, we have
@@
[…]
@@
We may assume that $M\ll N$. Following the proof of [13, Theorem 1.8], we construct a system of affine-linear forms $\Psi:\mathbb{Z}^{t-1}\to\mathbb{Z}^{t}$ of complexity $\leq 1$ and size $\lVert\Psi\rVert\ll 1$, such that $\Psi$ is injective with image exactly the full affine sublattice of $\mathbb{Z}^{t}$ defined by $\mathbf{c}\cdot\mathbf{x}=M$. Indeed, we can find an element $\mathbf{y}$ of this affine sublattice of size $\left|\mathbf{y}\right|\ll N$, and a $t\times(t-1)$-matrix $A$ of rank $t-1$ with entries in $\mathbb{Z}$ bounded in absolute value by $O(1)$, such that $A\cdot\mathbb{Z}^{t-1}$ is the kernel of $\mathbf{x}\mapsto\mathbf{c}\cdot\mathbf{x}$. Then the affine-linear system
@@
does what we want. Let us compare the main term in Theorem 1.4 for this system $\Psi$ and the convex set $X=\Psi^{-1}([0,N]^{t})$ to the above. Let $V\subseteq\mathbb{R}^{t}$ be the affine subspace defined by $\mathbf{c}\cdot\mathbf{x}=M$. The $(n-1)$-form in the definition of $\tau_{\infty}$, let us call it $\omega$, was chosen in the proof of [23, Proposition 25] so that $V\cap\mathbb{Z}^{t}$ has covolume $1$ in $V$. Hence, $\int_{[0,1)^{t-1}}\Psi^{*}(\omega)=\int_{\Psi([0,1)^{t-1})}\omega=1$, and therefore $\tau_{\infty}=\int_{\Psi(X)}\omega=\operatorname{vol}(X)$.
@@ 11.2. Three primes theorem with prescribed primitive roots
shows that $\tilde{\sigma}_{\mathbf{a},N}(q)=\sigma_{\mathbf{a},\Psi}(q)$ for all $q\in\mathbb{N}$. For even $N$, this quantity is clearly equal to 00 whenever $2\mid q$, hence, we may assume that $N$ is odd. The only significant difference between (11.3) and the asymptotic from Theorem 1.2 is that between $\tilde{\mathscr{D}}_{\mathbf{a}}$ and $\mathscr{D}_{\mathbf{a}}$. Let $m:=\min\{\nu_{2}(a_{i}):1\leq i\leq 3\}$ and $M:=\max\{\nu_{2}(a_{i}):1\leq i\leq 3\}$, so $\tilde{\mathscr{D}}_{\mathbf{a}}=2^{m}q$ and $\mathscr{D}_{\mathbf{a}}=2^{M}q$ for some odd $q$. We need to show that
@@
[…]
$\delta(a,b,2^{l}q)=\begin{cases}0&\text{ if }b\equiv 0\bmod 2,\\ 2^{-(l-1)}\delta(a,b,q)\left(1-\left(\frac{\Delta_{a}}{b}\right)\right)&\text{if }b\not\equiv 0\bmod 2\text{ and }1\leq v_{2}(a)\leq l,\\ 2^{-(l-1)}\delta(a,b,q)&\text{ if }b\not\equiv 0\bmod 2\text{ otherwise.}\end{cases}$
@@
$\sum_{\begin{subarray}{c}x_{i}\bmod 2^{l}q\\ x_{i}\equiv c_{i}\bmod q\\ x_{i}\equiv 1\bmod 2\end{subarray}}\left(\frac{\Delta_{a_{i}}}{x_{i}}\right)=\sum_{\begin{subarray}{c}y_{i}\bmod\Delta_{a_{i}}\\ y_{i}\equiv b_{i}\bmod(2q,\Delta_{a_{i}})\end{subarray}}\left(\frac{\Delta_{a_{i}}}{x_{i}}\right)\sum_{\begin{subarray}{c}x_{i}\bmod 2^{l}q\\ x_{i}\equiv y_{i}\bmod\Delta_{a_{i}}\\ x_{i}\equiv b_{i}\bmod 2q\end{subarray}}1=\frac{2^{l}q}{\operatorname{lcm}(\Delta_{a_{i}},2q)}\sum_{\begin{subarray}{c}y_{i}\bmod\Delta_{a_{i}}\\ y_{i}\equiv b_{i}\bmod(2q,\Delta_{a_{i}})\end{subarray}}\left(\frac{\Delta_{a_{i}}}{x_{i}}\right)=0.$
@@ 11.3. Three-term arithmetic progressions in primes with prescribed primitive roots
which is true for $(p-1)(p-2)$ vectors $(n_{1},n_{2})\in(\mathbb{Z}/p\mathbb{Z})^{2}$. For such $(n_{1},n_{2})$, we consider the congruences
@@
    There are $3(p-3)$ pairs $(n_{1},n_{2})$ satisfying (11.10) for which exactly one of the congruences is satisfied. Indeed, if this congruence is (11.11), then $n_{1}\equiv 1\bmod p$ and $n_{2}\not\equiv 0\bmod p$. Together with (11.10), this leaves $p-3$ choices for $n_{2}$. The other cases are similar.
@@
Table 11.1 below summarizes the value of the remaining part of $\mathfrak{S}(\mathbf{a},\Psi)$ for some choices of $\mathbf{a}$. Note that in case $\mathbf{a}=(2,3,6)$, even though the main term is zero, there are integers $(n_{1},n_{2})$ such that $(n_{1},n_{1}+n_{2},n_{1}+2n_{2})$ are primes with primitive roots $2,3,6$, respectively, such as $(3,4),(3,28),(3,40)$. Inserting $n_{1}=3$ and taking $p=3+n_{2}$ turns $\Psi$ into the system $(p,2p-3)$ of infinite complexity, about which we can not say anything. In particular, even under GRH we don’t know whether there are infinitely many pairs $(n_{1},n_{2})$ as above.
@@
$2$ $2$ $3$ $\frac{42}{25}\mathscr{A}_{2}^{2}\mathscr{A}_{3}$
$3$ $3$ $3$ $\frac{84}{25}\mathscr{A}_{3}^{3}$
$2$ $2$ $5$ $\frac{13500}{6859}\mathscr{A}_{2}^{2}\mathscr{A}_{5}$
$3$ $3$ $5$ $\frac{11340}{6859}\mathscr{A}_{3}^{2}\mathscr{A}_{5}$
$5$ $5$ $5$ $\frac{10000}{6859}\mathscr{A}_{5}^{3}$
$2$ $3$ $5$ $\frac{11340}{6859}\mathscr{A}_{2}\mathscr{A}_{3}\mathscr{A}_{5}$
$2$ $2$ $6$ $\frac{42}{25}\mathscr{A}_{2}^{2}\mathscr{A}_{6}$
$2$ $3$ $6$ 00
$6$ $6$ $6$ $\frac{42}{25}\mathscr{A}_{6}^{3}$
$2$ $5$ $6$ $\frac{11760}{6859}\mathscr{A}_{2}\mathscr{A}_{5}\mathscr{A}_{6}$
@@
$7$ $7$ $7$ $\frac{161084}{68921}\mathscr{A}_{7}^{3}$
$2$ $2$ $7$ $\frac{134050}{68921}\mathscr{A}_{2}^{2}\mathscr{A}_{7}$
$2$ $3$ $7$ $\frac{112602}{68921}\mathscr{A}_{2}\mathscr{A}_{3}\mathscr{A}_{7}$
$2$ $5$ $7$ $\frac{938350000}{472729139}\mathscr{A}_{2}\mathscr{A}_{5}\mathscr{A}_{7}$
$3$ $3$ $7$ $\frac{570192}{344605}\mathscr{A}_{3}^{2}\mathscr{A}_{7}$
@@
$2$ $3$ $10$ $\frac{54138}{34295}\mathscr{A}_{2}\mathscr{A}_{3}\mathscr{A}_{10}$
$2$ $5$ $10$ $\frac{14000}{6859}\mathscr{A}_{2}\mathscr{A}_{5}\mathscr{A}_{10}$
$5$ $5$ $10$ $\frac{1000}{361}\mathscr{A}_{5}^{2}\mathscr{A}_{7}$
@@
holds for $(p-2)(p-3)+p-1$ values of $n$. As before, we need to discuss the congruences
@@
    For exactly $4(p-3)$ values of $n$ satisfying (11.14), exactly one of the congruences is satisfied. More concretely, we obtain $p-3$ cases if we suppose (11.15), $p-2$ cases for (11.16), $p-3$ cases for (11.17), and $p-4$ cases for (11.18).
@@
[…]
@@
$2$ $2$ $2$ $3$ $\frac{72576}{78125}\mathscr{A}_{2}^{3}\mathscr{A}_{3}$
$2$ $2$ $2$ $5$ $\frac{252000}{130321}\mathscr{A}_{2}^{3}\mathscr{A}_{5}$
$2$ $2$ $3$ $3$ $\frac{63504}{15625}\mathscr{A}_{2}^{2}\mathscr{A}_{3}^{2}$
$2$ $2$ $3$ $5$ $\frac{508032}{130321}\mathscr{A}_{2}^{2}\mathscr{A}_{3}\mathscr{A}_{5}$
$2$ $2$ $3$ $6$ 00
$2$ $2$ $5$ $5$ $\frac{604800}{130321}\mathscr{A}_{2}^{2}\mathscr{A}_{5}^{2}$
$2$ $2$ $5$ $6$ $\frac{2159136}{651605}\mathscr{A}_{2}^{2}\mathscr{A}_{5}\mathscr{A}_{6}$
$2$ $2$ $6$ $6$ $\frac{72576}{78125}\mathscr{A}_{2}^{2}\mathscr{A}_{6}^{2}$
$2$ $3$ $3$ $3$ 00
$2$ $3$ $3$ $5$ $\frac{508032}{130321}\mathscr{A}_{2}\mathscr{A}_{3}^{2}\mathscr{A}_{5}$
@@
$2$ $3$ $3$ $6$ 00
$2$ $3$ $6$ $6$ $\frac{63504}{15625}\mathscr{A}_{2}\mathscr{A}_{3}\mathscr{A}_{6}^{2}$
$2$ $5$ $5$ $5$ $\frac{403200}{130321}\mathscr{A}_{2}\mathscr{A}_{5}^{3}$
$2$ $5$ $5$ $6$ $\frac{2286144}{651605}\mathscr{A}_{2}\mathscr{A}_{5}^{2}\mathscr{A}_{6}$
$2$ $6$ $6$ $6$ $\frac{244944}{78125}\mathscr{A}_{2}\mathscr{A}_{6}^{3}$
$3$ $3$ $3$ $5$ 00
$3$ $3$ $3$ $6$ 00
$3$ $3$ $5$ $5$ $\frac{3048192}{651605}\mathscr{A}_{3}^{2}\mathscr{A}_{5}^{2}$
$3$ $3$ $5$ $6$ $\frac{2159136}{651605}\mathscr{A}_{3}^{2}\mathscr{A}_{5}\mathscr{A}_{6}$
$3$ $3$ $6$ $6$ $\frac{63504}{15625}\mathscr{A}_{3}^{2}\mathscr{A}_{6}^{2}$
@@
    For each $p\in M_{2}$, we have $[K/\mathbb{Q},p]=C_{p}$.
This result answers a question concerning the existence of locally cyclic $\operatorname{S}_{n}$-extensions posed in [8, p.479] and generalises the case $n=5$ proved in [24, Theorem 5.5]. The proof follows the same strategy, except that we need to replace an application of the Green-Tao-Ziegler theorem in the case $n=5$ by Corollary 1.6, as for larger $n$ it becomes necessary to deal with splitting conditions in nonabelian fields. This yields the following new ingredient in the proof of Theorem A.1, which in case $R=1$ gives a positive answer to the question of Kim and König mentioned in the introduction of the main part. For $k\in\mathbb{N}$, we denote the splitting field of the polynomial
@@
[…]
@@
Then $q_{1},q_{2},q_{3}$ are pairwise coprime, $n$ divides $q_{1}$ and $n-1$ divides $q_{2}$. In particular, $n$ is invertible modulo $q_{2}q_{3}$ and $n-1$ is invertible modulo $q_{1}q_{3}$. As either $n$ or $n-1$ is even, we moreover see that $2\nmid q_{3}$. By the Chinese remainder theorem, we find $(s,t)\in(\mathbb{Z}/\mathscr{D}\mathbb{Z})^{2}$ that satisfy
@@
We have verified all hypotheses of Corollary 1.6. Hence, there are infinitely many pairs $(s,t)\in X\cap\mathbb{Z}^{2}$ for which $\psi_{1}(s,t),\psi_{2}(s,t),\psi_{3}(s,t)$ are pairwise distinct primes with the prescribed Artin symbols. Now the condition $[K_{1}/\mathbb{Q},\psi_{1}(s,t)]=C_{1}$ is equivalent to $\psi_{1}(s,t)\equiv u\bmod Rn$. For $(s,t)\in X$, we have $\psi_{1}(s,t)=(-1)^{n-1}t=|t|$, so these conditions translate to $|t|\equiv 1\bmod n$ and $t\equiv k\bmod R$. Similarly, the condition $[K_{1}/\mathbb{Q},\psi_{2}(s,t)]=C_{2}$ translates to $|s|\equiv 1\bmod n-1$ and $s\equiv l\bmod R$, and the condition $[K_{3}/\mathbb{Q},\psi_{3}(s,t)]=C_{3}$ means exactly that $|(n-1)^{n-1}t-n^{n}s|=(n-1)^{n-1}t-n^{n}s$ splits completely in $F_{n-1}$. ∎

### v3

@@ 9.4. W-trick
The following lemma is a simple reduction completely analogous to the deduction of [14, Theorem 5.1] from [14, Theorem 5.2].
@@
We need the following estimate in our application of the $W$-trick in Lemma 9.9. Its proof is, essentially, contained in the proof of [14, Lemma 1.3].
@@
so that $\psi_{i}(Wn+c)=W\psi_{i,c}(n)+b_{i}(c)$. Note that if $\gcd(W,\psi_{i}(c))>1$, then $\gcd(\psi_{i}(Wn+c),W)>1$ for any $Wn+c\in X$. As $W\ll_{\mathbf{a}}(\log N)^{2/3}$ and $\psi_{i}\geq N^{8/10}$ on $X$, this shows that $\psi_{i}(Wn+c)$ can not be prime if $N$ is large enough. Hence, identifying $W]^{s}$ with $(\mathbb{Z}/W\mathbb{Z})^{s}$ in the obvious way, we see that
@@
In [14, §4], the main theorem of [14] is first reduced to [14, Theorem 4.1], removing the archimedean factor, and then futher to [14, Theorem 4.5], extending $\Psi$ to a system in normal form.
@@
As we have a pseudorandom majorant already, the deduction of Theorem 1.4 from Proposition 2.1 follows the same steps as in §§9.39.5 with the obvious modifications. For example, we rely on the results concerning $\tau_{\mathbf{K},\mathbf{C},\Psi}(q)$ established in §3.2 instead of the analogous properties of $\sigma_{\mathbf{a},\Psi}(q)$ established in §3.3. Again, the situation is somewhat simpler, as the function $w(\cdot)$ is fixed once and for all and does not depend on any of the other parameters. The argument is essentially the same as in [14, §§4-7].
@@ 11.1. One linear equation in primes with prescribed Artin symbols
Here we show that the main term Kane’s result [24, Theorem 2] for $t\geq 3$ matches up with a special case or our Theorem 1.2. We may formulate Kane’s asymptotic as follows: for $t\geq 3$ and $\mathbf{K},\mathbf{C}$ as in Theorem 1.4, $\mathbf{c}=(c_{1},\ldots,c_{t})\in\mathbb{Z}^{t}$ with $\gcd(c_{1},\ldots,c_{t})=1$ and $N,M\in\mathbb{N}$, we have
@@
where $\Delta_{\mathbf{K}}=\operatorname{lcm}(|\Delta_{K_{i}^{\operatorname{ab}}}|\ :\ 1\leq i\leq t)$ is the least common multiple of the discriminants of the maximal abelian subfields $K_{i}^{\operatorname{ab}}$ of $K_{i}$ and
$\tau_{\infty}$ $=\int_{\begin{subarray}{c}x_{i}\in[0,N]\\ \sum_{i}c_{i}x_{i}=M\end{subarray}}\left(\sum_{i=1}^{t}\frac{c_{i}}{\lVert\mathbf{c}\rVert_{2}^{2}}\frac{\partial}{\partial x_{i}}\right)dx_{1}\wedge dx_{2}\wedge\ldots\wedge dx_{t},$
@@
We may assume that $M\ll N$. Following the proof of [14, Theorem 1.8], we construct a system of affine-linear forms $\Psi:\mathbb{Z}^{t-1}\to\mathbb{Z}^{t}$ of complexity $\leq 1$ and size $\lVert\Psi\rVert\ll 1$, such that $\Psi$ is injective with image exactly the full affine sublattice of $\mathbb{Z}^{t}$ defined by $\mathbf{c}\cdot\mathbf{x}=M$. Indeed, we can find an element $\mathbf{y}$ of this affine sublattice of size $\left|\mathbf{y}\right|\ll N$, and a $t\times(t-1)$-matrix $A$ of rank t1t-1 with entries in $\mathbb{Z}$ bounded in absolute value by $O(1)$, such that $A\cdot\mathbb{Z}^{t-1}$ is the kernel of $\mathbf{x}\mapsto\mathbf{c}\cdot\mathbf{x}$. Then the affine-linear system
@@
does what we want. Let us compare the main term in Theorem 1.4 for this system $\Psi$ and the convex set $X=\Psi^{-1}([0,N]^{t})$ to the above. Let $V\subseteq\mathbb{R}^{t}$ be the affine subspace defined by $\mathbf{c}\cdot\mathbf{x}=M$. The (n1)(n-1)-form in the definition of $\tau_{\infty}$, let us call it $\omega$, was chosen in the proof of [24, Proposition 25] so that $V\cap\mathbb{Z}^{t}$ has covolume $1$ in $V$. Hence, $\int_{[0,1)^{t-1}}\Psi^{*}(\omega)=\int_{\Psi([0,1)^{t-1})}\omega=1$, and therefore $\tau_{\infty}=\int_{\Psi(X)}\omega=\operatorname{vol}(X)$.
@@ 11.2. Three primes theorem with prescribed primitive roots
shows that $\tilde{\sigma}_{\mathbf{a},N}(q)=\sigma_{\mathbf{a},\Psi}(q)$ for all $q\in\mathbb{N}$. For even $N$, this quantity is clearly equal to 0 whenever $2\mid q$, hence, we may assume that $N$ is odd. The only significant difference between (11.3) and the asymptotic from Theorem 1.2 is that between $\tilde{\mathscr{D}}_{\mathbf{a}}$ and $\mathscr{D}_{\mathbf{a}}$. Let $m:=\min\{\nu_{2}(a_{i}):1\leq i\leq 3\}$ and $M:=\max\{\nu_{2}(a_{i}):1\leq i\leq 3\}$, so $\tilde{\mathscr{D}}_{\mathbf{a}}=2^{m}q$ and $\mathscr{D}_{\mathbf{a}}=2^{M}q$ for some odd $q$. We need to show that
$\tilde{\sigma}_{\mathbf{a},N}(2^{M}q)$ $=\tilde{\sigma}_{\mathbf{a},N}(2^{m}q),$ $\text{ if }m\geq 2\text{ or }m=M,$ (11.4)
$\tilde{\sigma}_{\mathbf{a},N}(2^{M}q)$ $=\tilde{\sigma}_{\mathbf{a},N}(2)\tilde{\sigma}_{\mathbf{a},N}(q),$ $\text{ if }m=0\text{ and }M\geq 1.$ (11.5)
@@
$\sum_{\begin{subarray}{c}x_{i}\bmod 2^{l}q\\ x_{i}\equiv c_{i}\bmod q\\ x_{i}\equiv 1\bmod 2\end{subarray}}\left(\frac{\Delta_{a_{i}}}{x_{i}}\right)=\hskip-14.22636pt\sum_{\begin{subarray}{c}y_{i}\bmod\Delta_{a_{i}}\\ y_{i}\equiv b_{i}\bmod(2q,\Delta_{a_{i}})\end{subarray}}\hskip-14.22636pt\left(\frac{\Delta_{a_{i}}}{x_{i}}\right)\sum_{\begin{subarray}{c}x_{i}\bmod 2^{l}q\\ x_{i}\equiv y_{i}\bmod\Delta_{a_{i}}\\ x_{i}\equiv b_{i}\bmod 2q\end{subarray}}\hskip-14.22636pt1=\frac{2^{l}q}{\operatorname{lcm}(\Delta_{a_{i}},2q)}\hskip-14.22636pt\sum_{\begin{subarray}{c}y_{i}\bmod\Delta_{a_{i}}\\ y_{i}\equiv b_{i}\bmod(2q,\Delta_{a_{i}})\end{subarray}}\left(\frac{\Delta_{a_{i}}}{x_{i}}\right)=0.$
@@ 11.3. Three-term arithmetic progressions in primes with prescribed primitive roots
which is true for (p1)(p2)(p-1)(p-2) vectors $(n_{1},n_{2})\in(\mathbb{Z}/p\mathbb{Z})^{2}$. For such $(n_{1},n_{2})$, we consider the congruences
@@
    There are 3(p3)3(p-3) pairs $(n_{1},n_{2})$ satisfying (11.10) for which exactly one of the congruences is satisfied. Indeed, if this congruence is (11.11), then $n_{1}\equiv 1\bmod p$ and $n_{2}\not\equiv 0\bmod p$. Together with (11.10), this leaves p3p-3 choices for $n_{2}$. The other cases are similar.
@@
Table 11.1 below summarizes the value of the remaining part of $\mathfrak{S}(\mathbf{a},\Psi)$ for some choices of $\mathbf{a}$. Note that in case $\mathbf{a}=(2,3,6)$, even though the main term is zero, there are integers $(n_{1},n_{2})$ such that $(n_{1},n_{1}+n_{2},n_{1}+2n_{2})$ are primes with primitive roots $2,3,6$, respectively, such as $(3,4),(3,28),(3,40)$. Inserting $n_{1}=3$ and taking $p=3+n_{2}$ turns $\Psi$ into the system (p,2p3)(p,2p-3) of infinite complexity, about which we can not say anything. In particular, even under GRH we don’t know whether there are infinitely many pairs $(n_{1},n_{2})$ as above.
@@
22 22 33 $\frac{42}{25}\mathscr{A}_{2}^{2}\mathscr{A}_{3}$
33 33 33 $\frac{84}{25}\mathscr{A}_{3}^{3}$
22 22 55 $\frac{13500}{6859}\mathscr{A}_{2}^{2}\mathscr{A}_{5}$
33 33 55 $\frac{11340}{6859}\mathscr{A}_{3}^{2}\mathscr{A}_{5}$
55 55 55 $\frac{10000}{6859}\mathscr{A}_{5}^{3}$
22 33 55 $\frac{11340}{6859}\mathscr{A}_{2}\mathscr{A}_{3}\mathscr{A}_{5}$
22 22 66 $\frac{42}{25}\mathscr{A}_{2}^{2}\mathscr{A}_{6}$
22 33 66 0
66 66 66 $\frac{42}{25}\mathscr{A}_{6}^{3}$
22 55 66 $\frac{11760}{6859}\mathscr{A}_{2}\mathscr{A}_{5}\mathscr{A}_{6}$
@@
77 77 77 $\frac{161084}{68921}\mathscr{A}_{7}^{3}$
22 22 77 $\frac{134050}{68921}\mathscr{A}_{2}^{2}\mathscr{A}_{7}$
22 33 77 $\frac{112602}{68921}\mathscr{A}_{2}\mathscr{A}_{3}\mathscr{A}_{7}$
22 55 77 $\frac{938350000}{472729139}\mathscr{A}_{2}\mathscr{A}_{5}\mathscr{A}_{7}$
33 33 77 $\frac{570192}{344605}\mathscr{A}_{3}^{2}\mathscr{A}_{7}$
@@
22 33 $10$ $\frac{54138}{34295}\mathscr{A}_{2}\mathscr{A}_{3}\mathscr{A}_{10}$
22 55 $10$ $\frac{14000}{6859}\mathscr{A}_{2}\mathscr{A}_{5}\mathscr{A}_{10}$
55 55 $10$ $\frac{1000}{361}\mathscr{A}_{5}^{2}\mathscr{A}_{7}$
@@
holds for (p2)(p3)+p1(p-2)(p-3)+p-1 values of $n$. As before, we need to discuss the congruences
@@
    For exactly 4(p3)4(p-3) values of $n$ satisfying (11.14), exactly one of the congruences is satisfied. More concretely, we obtain p3p-3 cases if we suppose (11.15), p2p-2 cases for (11.16), p3p-3 cases for (11.17), and p4p-4 cases for (11.18).
@@ 77 $\frac{914838624}{353220125}\mathscr{A}_{7}^{4}$
Table 11.2. Value of $\mathscr{C}_{\mathbf{a}}$ for several values of $\mathbf{a}=(a,a,a,a)$
@@
22 22 22 33 $\frac{72576}{78125}\mathscr{A}_{2}^{3}\mathscr{A}_{3}$
22 22 22 55 $\frac{252000}{130321}\mathscr{A}_{2}^{3}\mathscr{A}_{5}$
22 22 33 33 $\frac{63504}{15625}\mathscr{A}_{2}^{2}\mathscr{A}_{3}^{2}$
22 22 33 55 $\frac{508032}{130321}\mathscr{A}_{2}^{2}\mathscr{A}_{3}\mathscr{A}_{5}$
22 22 33 66 0
22 22 55 55 $\frac{604800}{130321}\mathscr{A}_{2}^{2}\mathscr{A}_{5}^{2}$
22 22 55 66 $\frac{2159136}{651605}\mathscr{A}_{2}^{2}\mathscr{A}_{5}\mathscr{A}_{6}$
22 22 66 66 $\frac{72576}{78125}\mathscr{A}_{2}^{2}\mathscr{A}_{6}^{2}$
22 33 33 33 0
22 33 33 55 $\frac{508032}{130321}\mathscr{A}_{2}\mathscr{A}_{3}^{2}\mathscr{A}_{5}$
@@
22 33 33 66 0
22 33 66 66 $\frac{63504}{15625}\mathscr{A}_{2}\mathscr{A}_{3}\mathscr{A}_{6}^{2}$
22 55 55 55 $\frac{403200}{130321}\mathscr{A}_{2}\mathscr{A}_{5}^{3}$
22 55 55 66 $\frac{2286144}{651605}\mathscr{A}_{2}\mathscr{A}_{5}^{2}\mathscr{A}_{6}$
22 66 66 66 $\frac{244944}{78125}\mathscr{A}_{2}\mathscr{A}_{6}^{3}$
33 33 33 55 0
33 33 33 66 0
33 33 55 55 $\frac{3048192}{651605}\mathscr{A}_{3}^{2}\mathscr{A}_{5}^{2}$
33 33 55 66 $\frac{2159136}{651605}\mathscr{A}_{3}^{2}\mathscr{A}_{5}\mathscr{A}_{6}$
33 33 66 66 $\frac{63504}{15625}\mathscr{A}_{3}^{2}\mathscr{A}_{6}^{2}$
@@
    For each $p\in M_{2}$, we have $K/\mathbb{Q},p]=C_{p}$.
This result answers a question concerning the existence of locally cyclic $\operatorname{S}_{n}$-extensions posed in [8, p.479] and generalises the case $n=5$ proved in [25, Theorem 5.5]. The proof follows the same strategy, except that we need to replace an application of the theorem of Green and Tao in the case $n=5$ by Corollary 1.6, as for larger $n$ it becomes necessary to deal with splitting conditions in nonabelian fields. This yields the following new ingredient in the proof of Theorem A.1, which in case $R=1$ gives a positive answer to the question of Kim and König mentioned in the introduction of the main part. For $k\in\mathbb{N}$, we denote the splitting field of the polynomial
@@
$|t|$ $\text{ is a prime congruent to }1\bmod n,$ (A.2)
$|s|$ $\text{ is a prime congruent to }1\bmod n-1,$ (A.3)
$|(n-1)^{n-1}t-n^{n}s|$ $\text{ is a prime splitting completely in }F_{n-1}.$ (A.4)
@@
Then $q_{1},q_{2},q_{3}$ are pairwise coprime, $n$ divides $q_{1}$ and n1n-1 divides $q_{2}$. In particular, $n$ is invertible modulo $q_{2}q_{3}$ and n1n-1 is invertible modulo $q_{1}q_{3}$. As either $n$ or n1n-1 is even, we moreover see that $2\nmid q_{3}$. By the Chinese remainder theorem, we find $(s,t)\in(\mathbb{Z}/\mathscr{D}\mathbb{Z})^{2}$ that satisfy
@@
We have verified all hypotheses of Corollary 1.6. Hence, there are infinitely many pairs $(s,t)\in X\cap\mathbb{Z}^{2}$ for which $\psi_{1}(s,t),\psi_{2}(s,t),\psi_{3}(s,t)$ are pairwise distinct primes with the prescribed Artin symbols. Now the condition $K_{1}/\mathbb{Q},\psi_{1}(s,t)]=C_{1}$ is equivalent to $\psi_{1}(s,t)\equiv u\bmod Rn$. For $(s,t)\in X$, we have $\psi_{1}(s,t)=(-1)^{n-1}t=|t|$, so these conditions translate to $|t|\equiv 1\bmod n$ and $t\equiv k\bmod R$. Similarly, the condition $K_{1}/\mathbb{Q},\psi_{2}(s,t)]=C_{2}$ translates to $|s|\equiv 1\bmod n-1$ and $s\equiv l\bmod R$, and the condition $K_{3}/\mathbb{Q},\psi_{3}(s,t)]=C_{3}$ means exactly that $|(n-1)^{n-1}t-n^{n}s|=(n-1)^{n-1}t-n^{n}s$ splits completely in $F_{n-1}$. ∎